Ticket Number: HEA-NOC/20020510-02 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEANET <=> IRCSET
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020515 08:30 UTC +1
Ticket Closed: 20020516 08:48:43 UTC Problem End: 20020515 08:35 UTC +1
Upgrade of line from HEAnet to IRCSET from 256Kbps to 1Mbps.
Brief outage on line to IRCSET possible.
20020515 08:30 BN
Changes made on both IRCSET and HEAnet ends. No outage detected. Upgrade complete but ticked being kept open for monitoring.
20020516 09:48 BN
Line has remained stable. Ticket closed.
Time to Fix: